# Normalized solutions of one-dimensional defocusing NLS equations with nonlinear point interactions

### v2

@@ 1. Introduction
Let $p>2$ and $2<q<\min\left\{4,\frac{p}{2}+1\right\}$. There results:
@@
Observe that, when $(p,q)$ belongs to region A, the fact that $\mathcal{E}_{p,q}$ is constant on $\mu_{p,q},+\infty)$ is particularly interesting also from a variational point of view. Tipically, when NLS ground states at a certain mass do not exist, this is due to the presence of an energetically convenient problem at infinity that makes energy minimizing sequences converge weakly to zero, thus losing all their mass in the limit (the vanishing case in the concentration–compactness framework). On the contrary, this is not what happens here in region A. Indeed, since $\mathcal{E}_{p,q}(\mu)=\mathcal{E}_{p,q}(\mu_{p,q})$ for every $\mu>\mu_{p,q}$, but it is attained only when $\mu=\mu_{p,q}$, energy minimizing sequences with any mass $\mu$ strictly larger than the threshold will fail to be compact in $H_{\mu}^{1}(\mathbb{R})$ because only the partial amount of mass $\mu-\mu_{p,q}$ runs away at infinity (with vanishing energy), while the rest of the mass concentrates towards a non-zero weak limit that is a ground state with mass $\mu_{p,q}$.
@@
In this section we present the proof of the first result of the paper, namely Theorem 1.1. For the sake of simlicity, we split the proof in three preliminary lemmas. The first one concerns qualitative features of the solutions of (3).
@@
Combining the continuity of $u$ on $\mathbb{R}$ and $u^{\prime}$ on $\mathbb{R}\setminus\{0\}$, the boundary condition in (3) and (7), one sees that $u$ cannot vanish on $\mathbb{R}$ (otherwise, $u\equiv 0$ by the Cauchy–Lipschitz Theorem) and, consequently, that $\lambda\geq 0$ (otherwise $0<\left(\frac{p|\lambda|}{2}\right)^{\frac{1}{p-2}}\leq|u(x)|\to 0$, as $|x|\to 0$). Furthermore, owing again to the continuity of $u$ and $u^{\prime}$, the boundary condition in (3) and (7), there results that $u^{\prime}(0^{-})=-u^{\prime}(0^{+})\neq 0$. Thus, with a suitable change of variable, one finds that $u$ solves somehow the same Cauchy Problem on $\mathbb{R}^{+}$ and $\mathbb{R}^{-}$, which proves even symmetry. Finally, in order to obtain decreasing monotonicity on $\mathbb{R}^{+}$ of a positive solution $u$, assume by contradiction that $u^{\prime}(x)>0$ for some $x\in\mathbb{R}^{+}$. Then, (7) implies $u^{\prime}(y)=\sqrt{\lambda u(y)^{2}+\frac{2|u(y)|^{p}}{p}}>0$, for every $y\geq x$, contradicting $\lim_{s\to+\infty}u(s)=0$. ∎
@@
and set $t=t_{\lambda}(a)$ in (9) so that, in view of (8), there results
@@
Arguing as in the proof of Lemma 2.2, one cah check that $u\in H^{1}(\mathbb{R})$ is a positive solution of (3) with $\lambda=0$ if and only if $p<6$ and $u$ is of the form
@@
In this section we present the proofs of Theorems 1.21.3. Note that throughout the section we focus only on positive and even solutions of (2), strictly deacreasing on $\mathbb{R}^{+}$, since this is not restrictive by Lemma 2.1.
@@
It is sufficient to note that the map $\lambda\mapsto\mu(\lambda)$ defined in (18) is a bijection of $\mathbb{R}^{+}$ onto intself. ∎
@@
Let $p,q>2$ with $q\neq\frac{p}{2}+1$. There results
@@
Thus, expanding computations and multiplying both sides times $\frac{(6-p)[(2q-p-2)\overline{t}^{2}+p-2]}{p-2}$, which is negative for the chosen range of p,qp,\,q, there results
@@
When $p\in(2,6)$ and $q=4$, (3) admits a nonrivial solution in $H^{1}(\mathbb{R})$ for every $\lambda\geq 0$ again by Theorem 1.1(ii) and Theorem 1.1(iii)(a). Moreover, Lemma 3.2 gives
@@
Since we need it later, let us focus on the map that associates each $t\in(1,+\infty)$ with the unique positive solution $u_{t}\in H^{1}(\mathbb{R})$ of (3) with $\lambda>0$ obtained through (8)–(11), and $\infty$ with the solution $u_{\infty}\in H^{1}(\mathbb{R})$ of (3) with $\lambda=0$ given by (14) (note that this last abuse of notation is consistent since, setting $t=+\infty$ in (11), one obtains $\lambda=0$). One can check that this map is continuous from any interval $I\subseteq(1,+\infty]$ on which it is defined to $H^{1}(\mathbb{R})$. Indeed, by (8)–(11) and (14), one can seen that $u_{t}\to u_{\overline{t}}$ pointwise on $\mathbb{R}$ as $t\to\overline{t}\in I$. Moreover, since $(u_{t})_{t\in I}$ is equibounded in $H^{1}(\mathbb{R})$ and every $u_{t}$ is radially decreasing on $\mathbb{R}$ (by Lemma 2.1), we have $u_{t}\to u_{\overline{t}}$ in $L^{r}(\mathbb{R})$ for every $r\in(2,+\infty]$, whereas the convergence of $u_{t}$ to $u_{\overline{t}}$ in $L^{2}(\mathbb{R})$ is a direct consequence of the continuity of the map $\mu(t)$ defined by (19). Given that (11) also ensures that each value of $t$ is associated with a single $\lambda(t)$ and that the map $t\mapsto\lambda(t)$ is continuous, passing to the limit as $t\to\overline{t}$ in (3) and exploiting the convergences proved so far one can prove that $u_{t}\to u_{\overline{t}}$ in $H^{1}(\mathbb{R})$.
@@
Observe that, for every $p,q>2$, the map $\mu\mapsto\mathcal{E}_{p,q}(\mu)$ is non–positive and non–increasing on $(0,+\infty)$. To see that $\mathcal{E}_{p,q}(\mu)\leq 0$ for every $p,q>2$ and $\mu>0$ one can argue as follows. Let $(u_{n})_{n}\subset H_{\mu}^{1}(\mathbb{R})$ be such that $u_{n}\equiv c_{n}$ on $n^{2},n^{2}]$, $u_{n}\equiv 0$ on $\mathbb{R}\setminus[-n^{2}-1,n^{2}+1]$, and $u_{n}$ is linear both on $n^{2},n^{2}+1]$ and on $n^{2}-1,-n^{2}]$. Here, $c_{n}>0$ is chosen to guarantee that $\|u_{n}\|_{2}^{2}=\mu$. Since a direct computation easily shows that $c_{n}\sim n^{-1}\sqrt{\frac{\mu}{2}}$, as $n\to+\infty$, there results that $\mathcal{E}_{p,q}(\mu)\leq\lim_{n\to+\infty}E_{p,q}(u_{n})=0$. On the other hand, to see the monotonicity of $\mathcal{E}_{p,q}$, let $0<\mu_{1}<\mu_{2}$ and $(v_{n})_{n}\subset H_{\mu_{1}}^{1}(\mathbb{R})$ be such that $E_{p,q}(v_{n})\to\mathcal{E}_{p,q}(\mu_{1})$ as $n\to+\infty$. Without loss of generality, assume also that $(v_{n})_{n}\subset C_{0}^{\infty}(\mathbb{R})$. Set, now, $w_{n}(x):=v_{n}(x)+u_{n}(x-y_{n})$, where $(u_{n})_{n}\subset H_{\mu_{2}-\mu_{1}}^{1}(\mathbb{R})$ is defined as in the first part of the remark, but with mass $\mu_{2}-\mu_{1}$, and $y_{n}\in\mathbb{R}$ is such that the supports of $v_{n}$ and of $u_{n}(\cdot-y_{n})$ are disjoint for every $n$. Then, $(w_{n})_{n}\in H_{\mu_{2}}^{1}(\mathbb{R})$ and
@@
Let, then, $v_{t}=u+t\varphi$ for every $t>0$, for a fixed $\varphi\in H^{1}(\mathbb{R})$ such that $\varphi>0$. Clearly, $\mu_{t}:=\|v_{t}\|_{2}^{2}\downarrow\mu$ as $t\to 0^{+}$ and, setting $f(t):=E_{p,q}(v_{t})$ for every $t>0$, there results
@@
the last inequality coming again by lower semicontinuity. Since, by assumption, $\mathcal{E}_{p,q}(\overline{\mu})<\mathcal{E}_{p,q}(\mu)$ for every $\mu<\overline{\mu}$, there results $m=\overline{\mu}$. Hence, $u\in H_{\overline{\mu}}^{1}(\mathbb{R})$ and all the above inequalities are in fact equalities, so that $E_{p,q}(u)=\mathcal{E}_{p,q}(\overline{\mu})$, i.e. $u$ is a ground state at mass $\overline{\mu}$. ∎
@@
there results that $v\in H_{\mu}^{1}(\mathbb{R})$ and, since $q<\min\left\{4,\frac{p}{2}+1\right\}$ implies $0<\frac{q}{4-q}<\frac{p+2-q}{4-q}$, that $\mathcal{E}_{p,q}(\mu)\leq E_{p,q}(v)<0$ as soon as $\delta$ and $\mu$ are sufficiently close to zero. This proves that $\mathcal{E}_{p,q}(\mu)$ is strictly negative in a right neighborhood of 0, and thus for every $\mu>0$ by Remark 4.1.
@@
Conflict of interest The authors declare that they have no conflict of interest.

### v3

@@ 1. Introduction
Let $p>2$ and $2<q<\min\left\{4,\frac{p}{2}+1\right\}$. The following holds:
@@
Observe that, when $(p,q)$ belongs to region A, the fact that $\mathcal{E}_{p,q}$ is constant on $\mu_{p,q},+\infty)$ is particularly interesting also from a variational point of view. Typically, when NLS ground states at a certain mass do not exist, this is due to the presence of an energetically convenient problem at infinity that makes energy minimizing sequences converge weakly to zero, thus losing all their mass in the limit (the vanishing case in the concentration–compactness framework). On the contrary, this is not what happens here in region A. Indeed, since $\mathcal{E}_{p,q}(\mu)=\mathcal{E}_{p,q}(\mu_{p,q})$ for every $\mu>\mu_{p,q}$, but it is attained only when $\mu=\mu_{p,q}$, energy minimizing sequences with any mass $\mu$ strictly larger than the threshold will fail to be compact in $H_{\mu}^{1}(\mathbb{R})$ because only the partial amount of mass $\mu-\mu_{p,q}$ runs away at infinity (with vanishing energy), while the rest of the mass concentrates towards a non-zero weak limit that is a ground state with mass $\mu_{p,q}$.
@@
In this section we present the proof of the first result of the paper, namely Theorem 1.1. For the sake of simplicity, we split the proof in three preliminary lemmas. The first one concerns qualitative features of the solutions of (3).
@@
Combining the continuity of $u$ on $\mathbb{R}$ and $u^{\prime}$ on $\mathbb{R}\setminus\{0\}$, the boundary condition in (3) and (7), one sees that $u$ cannot vanish on $\mathbb{R}$ (otherwise, $u\equiv 0$ by the Cauchy–Lipschitz Theorem) and, consequently, that $\lambda\geq 0$ (otherwise $0<\left(\frac{p|\lambda|}{2}\right)^{\frac{1}{p-2}}\leq|u(x)|\to 0$, as $|x|\to+\infty$). Furthermore, owing again to the continuity of $u$ and $u^{\prime}$, the boundary condition in (3) and (7), it holds that $u^{\prime}(0^{-})=-u^{\prime}(0^{+})\neq 0$. Thus, with a suitable change of variable, one finds that $u$ solves somehow the same Cauchy Problem on $\mathbb{R}^{+}$ and $\mathbb{R}^{-}$, which proves even symmetry. Finally, in order to obtain decreasing monotonicity on $\mathbb{R}^{+}$ of a positive solution $u$, assume by contradiction that $u^{\prime}(x)>0$ for some $x\in\mathbb{R}^{+}$. Then, (7) implies $u^{\prime}(y)=\sqrt{\lambda u(y)^{2}+\frac{2|u(y)|^{p}}{p}}>0$, for every $y\geq x$, contradicting $\lim_{s\to+\infty}u(s)=0$. ∎
@@
and set $t=t_{\lambda}(a)$ in (9) so that, in view of (8), one obtains
@@
Arguing as in the proof of Lemma 2.2, one can check that $u\in H^{1}(\mathbb{R})$ is a positive solution of (3) with $\lambda=0$ if and only if $p<6$ and $u$ is of the form
@@
In this section we present the proofs of Theorems 1.21.3. Note that throughout the section we focus only on positive and even solutions of (2), strictly decreasing on $\mathbb{R}^{+}$, since this is not restrictive by Lemma 2.1.
@@
It is sufficient to note that the map $\lambda\mapsto\mu(\lambda)$ defined in (18) is a bijection of $\mathbb{R}^{+}$ onto itself. ∎
@@
Let $p,q>2$ with $q\neq\frac{p}{2}+1$. It holds that
@@
Thus, expanding computations and multiplying both sides times $\frac{(6-p)[(2q-p-2)\overline{t}^{2}+p-2]}{p-2}$, which is negative for the chosen range of p,qp,\,q, we have
@@
When $p\in(2,6)$ and $q=4$, (3) admits a nontrivial solution in $H^{1}(\mathbb{R})$ for every $\lambda\geq 0$ again by Theorem 1.1(ii) and Theorem 1.1(iii)(a). Moreover, Lemma 3.2 gives
@@
Since we need it later, let us focus on the map that associates each $t\in(1,+\infty)$ with the unique positive solution $u_{t}\in H^{1}(\mathbb{R})$ of (3) with $\lambda>0$ obtained through (8)–(11), and $\infty$ with the solution $u_{\infty}\in H^{1}(\mathbb{R})$ of (3) with $\lambda=0$ given by (14) (note that this last abuse of notation is consistent since, setting $t=+\infty$ in (11), one obtains $\lambda=0$). One can check that this map is continuous from any interval $I\subseteq(1,+\infty]$ on which it is defined to $H^{1}(\mathbb{R})$. Indeed, by (8)–(11) and (14), one can see that $u_{t}\to u_{\overline{t}}$ pointwise on $\mathbb{R}$ as $t\to\overline{t}\in I$. Moreover, since $(u_{t})_{t\in I}$ is equibounded in $H^{1}(\mathbb{R})$ and every $u_{t}$ is radially decreasing on $\mathbb{R}$ (by Lemma 2.1), we have $u_{t}\to u_{\overline{t}}$ in $L^{r}(\mathbb{R})$ for every $r\in(2,+\infty]$, whereas the convergence of $u_{t}$ to $u_{\overline{t}}$ in $L^{2}(\mathbb{R})$ is a direct consequence of the continuity of the map $\mu(t)$ defined by (19). Given that (11) also ensures that each value of $t$ is associated with a single $\lambda(t)$ and that the map $t\mapsto\lambda(t)$ is continuous, passing to the limit as $t\to\overline{t}$ in (3) and exploiting the convergences proved so far one can prove that $u_{t}\to u_{\overline{t}}$ in $H^{1}(\mathbb{R})$.
@@
Observe that, for every $p,q>2$, the map $\mu\mapsto\mathcal{E}_{p,q}(\mu)$ is non–positive and non–increasing on $(0,+\infty)$. To see that $\mathcal{E}_{p,q}(\mu)\leq 0$ for every $p,q>2$ and $\mu>0$ one can argue as follows. Let $(u_{n})_{n}\subset H_{\mu}^{1}(\mathbb{R})$ be such that $u_{n}\equiv c_{n}$ on $n^{2},n^{2}]$, $u_{n}\equiv 0$ on $\mathbb{R}\setminus[-n^{2}-1,n^{2}+1]$, and $u_{n}$ is linear both on $n^{2},n^{2}+1]$ and on $n^{2}-1,-n^{2}]$. Here, $c_{n}>0$ is chosen to guarantee that $\|u_{n}\|_{2}^{2}=\mu$. Since a direct computation easily shows that $c_{n}\sim n^{-1}\sqrt{\frac{\mu}{2}}$, as $n\to+\infty$, it follows that $\mathcal{E}_{p,q}(\mu)\leq\lim_{n\to+\infty}E_{p,q}(u_{n})=0$. On the other hand, to see the monotonicity of $\mathcal{E}_{p,q}$, let $0<\mu_{1}<\mu_{2}$ and $(v_{n})_{n}\subset H_{\mu_{1}}^{1}(\mathbb{R})$ be such that $E_{p,q}(v_{n})\to\mathcal{E}_{p,q}(\mu_{1})$ as $n\to+\infty$. Without loss of generality, assume also that $(v_{n})_{n}\subset C_{0}^{\infty}(\mathbb{R})$. Set, now, $w_{n}(x):=v_{n}(x)+u_{n}(x-y_{n})$, where $(u_{n})_{n}\subset H_{\mu_{2}-\mu_{1}}^{1}(\mathbb{R})$ is defined as in the first part of the remark, but with mass $\mu_{2}-\mu_{1}$, and $y_{n}\in\mathbb{R}$ is such that the supports of $v_{n}$ and of $u_{n}(\cdot-y_{n})$ are disjoint for every $n$. Then, $(w_{n})_{n}\in H_{\mu_{2}}^{1}(\mathbb{R})$ and
@@
Let, then, $v_{t}=u+t\varphi$ for every $t>0$, for a fixed $\varphi\in H^{1}(\mathbb{R})$ such that $\varphi>0$. Clearly, $\mu_{t}:=\|v_{t}\|_{2}^{2}\downarrow\mu$ as $t\to 0^{+}$ and, setting $f(t):=E_{p,q}(v_{t})$ for every $t>0$, one has
@@
the last inequality coming again by lower semicontinuity. Since, by assumption, $\mathcal{E}_{p,q}(\overline{\mu})<\mathcal{E}_{p,q}(\mu)$ for every $\mu<\overline{\mu}$, it follows that $m=\overline{\mu}$. Hence, $u\in H_{\overline{\mu}}^{1}(\mathbb{R})$ and all the above inequalities are in fact equalities, so that $E_{p,q}(u)=\mathcal{E}_{p,q}(\overline{\mu})$, i.e. $u$ is a ground state at mass $\overline{\mu}$. ∎
@@
it holds that $v\in H_{\mu}^{1}(\mathbb{R})$ and, since $q<\min\left\{4,\frac{p}{2}+1\right\}$ implies $0<\frac{q}{4-q}<\frac{p+2-q}{4-q}$, that $\mathcal{E}_{p,q}(\mu)\leq E_{p,q}(v)<0$ as soon as $\delta$ and $\mu$ are sufficiently close to zero. This proves that $\mathcal{E}_{p,q}(\mu)$ is strictly negative in a right neighborhood of 0, and thus for every $\mu>0$ by Remark 4.1.
@@
Conflict of interest. The authors declare that they have no conflict of interest.
Data availability statement. Data sharing not applicable to this article as no dataset was generated or analysed during the current study.